OpenStudy (anonymous):

[Warning Physics involved] The school bus with a total mass of 10 tonnes travels at a constant velocity of 40km/h down a long hill with a slope of 15 degrees. determine the braking force required to stop the bus within 50 metres from where the driver applies the brakes.

OpenStudy (anonymous):

come on where my physicians at ?? :/

OpenStudy (shamim):

we know F=ma F=

OpenStudy (shamim):

m=10000kg

OpenStudy (shamim):

a=?

OpenStudy (shamim):

v^2=u^2-2as

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

1/810 m/s^2????

OpenStudy (anonymous):

idk that seems kind of dodgy to me ...?

OpenStudy (anonymous):

yeah so we should square 3600 ?

OpenStudy (shamim):

initial velocity u=40km/h=40*1000/3600m/s

OpenStudy (shamim):

now we know v^2=u^2-2as 0^2=(100/9)^2-2a*50

OpenStudy (shamim):

now calculate acceleration a

OpenStudy (shamim):

now tell me the result for a

OpenStudy (shamim):

ok

OpenStudy (anonymous):

11.1

OpenStudy (shamim):

ok

OpenStudy (shamim):

force F=ma=10000*11.1

OpenStudy (anonymous):

111000? is that the final answer ?

OpenStudy (shamim):

it has another force which is a component of weight of ur bus

OpenStudy (shamim):

weight F=10000*10 and component of weight along the way is 10000*10*cos 75

OpenStudy (shamim):

final force is 111000-10000*10*cos 75

OpenStudy (shamim):

calculate plz

OpenStudy (anonymous):

111000-25881.9=85118.1